OpenStudy (anonymous):

Which equation could be used to solve this problem? The sum of two consecutive odd integers is 124. Find the integers. A. n + 3 = 124 B. 2n + 4 = 124 C. 2n + 2 = 124 D. 2n + 1 = 124

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@shamil98

OpenStudy (the_fizicx99):

Same as the other. 2n + 1 = 124 -1 -1 2n = 123 /2 /2

OpenStudy (anonymous):

so its D TOO

OpenStudy (anonymous):

@tHe_FiZiCx99 is it D to

OpenStudy (the_fizicx99):

Apparently